OpenStudy (anonymous):

I need serious help please!!!!!!!!!! I will give medal and be a fan if helped. F(x) =-16x^2 +22x+ 3 Part A: What are the x-intercepts of the graph of the f(x) ? Show your work Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work Part C: What are the steps you would use to graph f(x)? Make sure that you can use the answers obtained in Part A and Part B to draw the graph

terenzreignz (terenzreignz):

Part A. x-intercepts are going to be the values of x where the entire function is equal to zero.

OpenStudy (anonymous):

awesome thank you rare you able to do B as well?

terenzreignz (terenzreignz):

A isn't quite done yet, I've only told you how to do it.

OpenStudy (anonymous):

I figured so

terenzreignz (terenzreignz):

So... since you're ready to do B, I suppose you have your answers to A?

OpenStudy (anonymous):

haha sadly no I don't have the answer when I thought I did.

terenzreignz (terenzreignz):

Solve the quadratic equation \[\Large -16x^2+22x+3=0\]using whatever method you're comfortable with, maybe quadratic formula/

OpenStudy (anonymous):

can I use the quadratic formula

terenzreignz (terenzreignz):

Yup... go right ahead :)

OpenStudy (anonymous):

x=3/2,-1/8 is the answer right?

terenzreignz (terenzreignz):

Okay, that's right. Now, B is relatively straightforward... it's a maximum if the coefficient of x^2 is negative, and it's a minimum if the coefficient of x^2 is positive. The vertex of a quadratic equation of the form \[\Large \color{blue}ax^2 + \color{red}bx + \color{green}c\] always takes the form \[\Large \left(-\frac{\color{red}b}{2\color{blue}a}, \frac{4\color{blue}a\color{green}c - \color{red}b^2}{4\color{blue}a}\right)\] Just plug in the values and solve ^_^

OpenStudy (anonymous):

what Values do I plug in?

terenzreignz (terenzreignz):

a is the coefficient of the x^2 b is the coefficient of the x (the one with no exponent) c is the constant

terenzreignz (terenzreignz):

Anyway, you should already know these a, b, and c values, they are the same values as the one you used when you used the quadratic formula.

OpenStudy (anonymous):

sorry its just that i am in a bit of o hurry to get this assignment done because i am going hunting for the first time today so all I have been looking for today is straight up answers.

OpenStudy (anonymous):

and I forget what coefficient means and what u mean by constant

terenzreignz (terenzreignz):

I suggest you go hunting now, and enjoy, and then come back to this problem when you have nothing else in your mind, because doing maths like this tends to take focus, and patience. Also, I'll be suspended if I give answers straight out. :)

OpenStudy (anonymous):

where will u get suspended from?

terenzreignz (terenzreignz):

This website. And besides, if you didn't know what a, b, and c were, how did you use the quadratic formula? :/

OpenStudy (anonymous):

I used something to help me lets say for now.

terenzreignz (terenzreignz):

I see. Well, I'm trying to help you, but I need your cooperation and patience... but you certainly seem to be in a hurry, so just enjoy your hunting trip first, before relaxing and doing maths ^_^

OpenStudy (anonymous):

I apologize for lying to u

terenzreignz (terenzreignz):

No problem, just don't make a habit of it ^_^

OpenStudy (anonymous):

also I should tell you im home schooled and I have to get this done today or ill be screwed tomorrow and ill be out hunting until around 9pm

terenzreignz (terenzreignz):

Well then, put more effort into it, and try to put some speed, too. What are a, b, and c? (Also, no chance of reducing hunting time? Get your priorities straight)

OpenStudy (anonymous):

This has been planed for quit a while so it would have been messed up to reduce the time plus night time is when the hogs come out.

terenzreignz (terenzreignz):

Bummer. So I guess the best course of action is to just get this done as quickly as you can, and that begins with correctly identifying your values for a, b, and c. :)

OpenStudy (anonymous):

apparently so.

terenzreignz (terenzreignz):

So... figured them out yet? It's not nearly as hard as you might think, the values for a, b, and c are literally right in front of you :) Example: \[\large 2x^2 - 7x +9\] a here is 2, the number next to x^2 b here is -7, the number next to x c here is 9, the number with no x Now, try with your expression. \[\Large -16x^2+22x + 3\]

terenzreignz (terenzreignz):

Oh, and fair warning ^_^ I'm so sleepy right now, I could fall asleep anytime without remembering to notify or log out, so you might want to keep that in mind :)

OpenStudy (anonymous):

-16=a 22=b 3=c

terenzreignz (terenzreignz):

That's better. So plug into the vertex formula and find your vertex point ^_^

terenzreignz (terenzreignz):

Vertex Formula \[\Large \left(-\frac{\color{red}b}{2\color{blue}a}, \frac{4\color{blue}a\color{green}c - \color{red}b^2}{4\color{blue}a}\right)\]

OpenStudy (anonymous):

to be honest here I have never heard of the vortex point.

OpenStudy (anonymous):

Im I 9th grade FYI